Need help please
The perimeter of a rectangular photograph is 22 inches. The lenth of the photograph is 1 inch more than the width. What are the dimensions of the photograph?

Answer by scott8148(6628)

P = 2L + 2W ___ 22 = 2L + 2W ___ 11 = L + W

"The lenth of the photograph is 1 inch more than the width" ___ L = W + 1

substituting (for the value of L) ___ 11 = (W + 1) + W ___ 11 = 2W + 1 ___ 5 = W

substituting ___ L = (5) + 1 = 6